The points (-1,3) and (0,2) are on a curve, and at any point (x,y) on the curve 𝑑^2𝑦/𝑑π‘₯^2 = 2 βˆ’ 4π‘₯. Find an equation of the curve.


1
Expert's answer
2021-03-01T07:10:54-0500

To find the equation of the curve, we first find the general solution of the given differential equation.



"\dfrac{d^2y}{dx^2} = 2-4x"

We can rewrite the above as:


"d^2y =(2-4x)dx^2\\
d(dy) = [(2-4x)dx]dx"

Integrating the above:


"\int d(dy) = \int[(2-4x)dx]dx\\
dy = (2x - 2x^2 +c_1)dx"

Integrating again:


"\int dy = \int(2x - 2x^2 +c_1)dx\\
\bold{y =x^2-\frac{2x^3}{3}+c_1x + c_2}"

Which is the general solution of the curve.


At point (-1,3), where x = -1 and y = 3:


"3 = (-1)^2 - \frac{2(-1)^3}{3} +c_1(-1) + c_2\\
3 = 1 + \frac{2}{3} - c_1+c_2\\
3-1-\frac{2}{3} = -c_1+c_2\\
\frac{4}{3} = -c_1+c_2 \cdots \cdots (eqn 1)"

At point (0,2), where x = 0 and y=2;

"2 = (0)^2 - \frac{2(0)^3}{3} +c_1(0) + c_2\\
2 = c_2 \cdots \cdots (eqn 2)"

Substitute eqn 2 into eqn 1:


"\frac{4}{3} = -c_1+2\\
c_1 = 2 - \frac{4}{3}\\
c_1 = \frac{2}{3}"

Thus the equation of the curve is:


"\bold{y =x^2-\frac{2x^3}{3}+\frac{2x}{3} + 2}"

OR


"\bold{y =2+\frac{2x}{3}+x^2-\frac{2x^3}{3} }"